OpenStudy (anonymous):

Need to find X and Y x=3y+7 x=2y-1 please explain how you came up with your answer

OpenStudy (anonymous):

You make the assumption that the x from equation 1 is the same as the x from equation 2 somewhere in the plane, and so you set them equal to each other but write each x in terms of y: 3y+7=2y-1 3y-2y=-1-7 y=-8 after solving for y, you find x by substituting -8 for y in either of the two original equations (I choose the second equation): x=2(-8)-1=-16-1=-17 the geometric interpretation of a unique solution of a system of equations is the point of intersection between the two lines, so the solution set can be written as {(-17,-8)}